# Generalization Error of the Tilted Empirical Risk

### v2

@@ Appendix G Proofs and details of Section 6
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\alpha\frac{(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}.\end{split}$ (104)
Using Theorem 3.13, the final result follows. ∎
@@
Proposition G.1 (Gibbs posterior).
@@
with $\hat{\mathrm{R}}(H,S)$ the linear empirical risk (1), is the Gibbs posterior,
@@
where $\gamma=\frac{1}{\alpha}-1$. As shown by Dvijotham and Todorov (2012), the solution to (106) is the Gibbs posterior,
@@
Proposition G.2.
Under Assumption 3.1 when training with the Gibbs posterior, (28), the following upper bound holds on the expected tilted generalization error,
@@
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&=\mathrm{R}(H,P_{H}\otimes\mu)-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]\\ &\quad+\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}).\end{split}$ (108)
@@
$\begin{split}&\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\gamma}\log\big{(}\exp(\gamma\ell(H,Z_{i}))\big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{)}\\ &\leq\frac{-\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}\exp(-2\gamma M).\end{split}$
@@
From Proposition G.2, we can observe that for $\gamma>0$ the upper bound on the tilted generalization error under the Gibbs posterior is tighter than the upper bound on the generalization error of the Gibbs posterior under linear empirical risk given in Aminian et al. (2021a).
@@
Appendix H Other Bounds
In this section, we provide upper bounds via Rademacher complexity, stability and PAC-Bayesian approaches. The results are based on the assumption of bounded loss functions (Assumption 3.1).
H.1 Rademacher Complexity
Inspired by the work (Bartlett and Mendelson, 2002), we provide an upper bound on the tilted generalization error via Rademacher complexity analysis. For this purpose, we need to define the Rademacher complexity.
As in Bartlett and Mendelson (2002), for a hypothesis set $\mathcal{H}$ of functions $h:\mathcal{X}\mapsto\mathcal{Y}$, the Rademacher complexity with respect to the dataset $S$ is
@@ H.1 Rademacher Complexity
Given Assumptions 3.1 and assuming the loss function is $M_{\ell^{\prime}}$-Lipschitz-continuous in a binary classification problem, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\widehat{\mathrm{gen}}_{\gamma}(h,S)\leq 2\exp(|\gamma|M)M_{\ell^{\prime}}\hat{\mathfrak{R}}_{\mathrm{S}}(\mathcal{H})+\frac{3(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}}.\end{split}$
@@
where $\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})$ can be bounded using Proposition 3.2. The second term can be bounded by using Proposition H.1 and Massart’s lemma (Lemma C.13). ∎
Similar to Remark 3.8, assuming $\gamma=O(1/\sqrt{n})$, we have the convergence rate of $O(1/\sqrt{n})$ for the tilted generalization error. For an infinite hypothesis space, covering number bounds can be applied to the empirical Rademacher complexity, see, e.g., (Kakade et al., 2008). We note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms.
H.2 A Stability Bound
In this section, we also study the upper bound on the tilted generalization error from the stability perspective (Bousquet and Elisseeff, 2002b). In the stability approach, (Bousquet and Elisseeff, 2002b), the learning algorithm is a deterministic function of $S$.
@@ H.2 A Stability Bound
Under Assumption 3.1, the following upper bound holds with probability at least $(1-\delta)$ under distribution $P_{S}$, $\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\tilde{Z}}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$ (113)
@@
We also consider the uniform stability as in Bousquet and Elisseeff (2002b).
@@ H.3 A PAC-Bayesian Bound
Under Assumption 3.1, the following upper bound holds on the conditional expected tilted generalization error with probability at least $(1-\delta)$ under the distribution $P_{S}$; for any $\eta>0$, $\begin{split}&|\mathbb{E}_{\rho_{h}}[\mathrm{gen}_{\gamma}(H,S)]|\leq\frac{\max(1,\exp(-2\gamma M))(1-\exp(\gamma M))^{2}}{8|\gamma|}\\ &\qquad+\frac{L\eta A^{2}}{8n}+\frac{L(\mathrm{KL}(\rho_{h}\|Q_{h})+\log(1/\delta))}{\eta},\end{split}$ (116) where $Q_{h}$ and $\rho_{h}$ are prior and posterior distributions over the hypothesis space, respectively.

### v3

@@ H.3 The KL-Regularized TERM Problem
$\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\leq\alpha\frac{(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}.$ (114)
Using Theorem H.11, the final result follows. ∎
Similar to Corollary H.13, we derive the following upper bound on the absolute value of the expected tilted generalization error of the tilted generalization error.
Corollary H.16.
Under the same assumptions in Theorem H.15, the absolute value of the expected tilted generalization error of the tilted Gibbs posterior satisfies
$|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.$ (115)
Remark H.17 (Convergence rate).
Remark H.18 (Discussion of $\gamma$).
From the upper bound in Theorem H.15, we can observe that under $\gamma\rightarrow 0$ and Assumption H.1, the upper bound converges to the upper bound on the Gibbs posterior (Aminian et al., 2021a). For positive tilt $(\gamma>0)$, and sufficient large value of $n$, the upper bound in Theorem H.15, can be tighter than the upper bound on the Gibbs posterior.
@@
Proposition H.19 (Gibbs posterior).
@@
with $\hat{\mathrm{R}}(H,S)$ the linear empirical risk (1), and the Gibbs posterior,
@@
where $\gamma=\frac{1}{\alpha}-1$. As shown by Dvijotham & Todorov (2012), the solution to (117) is the Gibbs posterior,
@@
Proposition H.20.
Under Assumption H.1 when training with the Gibbs posterior, (26), the following upper bound holds on the expected tilted generalization error,
@@
$\overline{\mathrm{gen}}_{\gamma}(H,S)=\mathrm{R}(H,P_{H}\otimes\mu)-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]+\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\ell(H,Z_{i})]-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}).$ (119)
@@
$\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\gamma}\log\big{(}\exp(\gamma\ell(H,Z_{i}))\big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{)}\leq\frac{-\mathrm{Var}(\exp(\gamma\ell(H,Z)))}{2\gamma}\exp(-2\gamma M).$
@@
H.4 Other Bounds
In this section, we provide upper bounds via Rademacher complexity (Bartlett & Mendelson, 2002) and PAC-Bayesian approaches (Alquier, 2021) under bounded loss functions assumption. The results are based on the assumption of bounded loss functions (Assumption H.1).
H.4.1 Rademacher Complexity
Inspired by the work (Bartlett & Mendelson, 2002), we provide an upper bound on the tilted generalization error via Rademacher complexity analysis. For this purpose, we need to define the Rademacher complexity.
As in Bartlett & Mendelson (2002), for a hypothesis set $\mathcal{H}$ of functions $h:\mathcal{X}\mapsto\mathcal{Y}$, the Rademacher complexity with respect to the dataset $S$ is
@@ H.4.1 Rademacher Complexity
Given Assumptions H.1 and assuming the loss function is $M_{\ell^{\prime}}$-Lipschitz-continuous in a binary classification problem, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\widehat{\mathrm{gen}}_{\gamma}(h,S)\leq 2\exp(|\gamma|M)M_{\ell^{\prime}}\hat{\mathfrak{R}}_{\mathrm{S}}(\mathcal{H})+\frac{3(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}}.\end{split}$
@@
where $\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})$ can be bounded using Proposition H.2. The second term can be bounded by using Proposition H.21 and Massart’s lemma (Lemma C.13). ∎
Similar to Remark H.7, assuming $\gamma=O(1/\sqrt{n})$, we have the convergence rate of $O(1/\sqrt{n})$ for the tilted generalization error. For an infinite hypothesis space, covering number bounds can be applied to the empirical Rademacher complexity, see, e.g., (Kakade et al., 2008). We note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms.
H.4.2 A Stability Bound
In this section, we also study the upper bound on the tilted generalization error from the stability perspective (Bousquet & Elisseeff, 2002b). In the stability approach, (Bousquet & Elisseeff, 2002b), the learning algorithm is a deterministic function of $S$.
@@ H.4.2 A Stability Bound
Under Assumption H.1, the following upper bound holds with probability at least $(1-\delta)$ under distribution $P_{S}$, $\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\tilde{Z}}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$ (124)
@@
We also consider the uniform stability as in (Bousquet & Elisseeff, 2002b).
@@ H.4.3 A PAC-Bayesian Bound
Under Assumption H.1, the following upper bound holds on the conditional expected tilted generalization error with probability at least $(1-\delta)$ under the distribution $P_{S}$; for any $\eta>0$, $\begin{split}&|\mathbb{E}_{\rho_{h}}[\mathrm{gen}_{\gamma}(H,S)]|\leq\frac{\max(1,\exp(-2\gamma M))(1-\exp(\gamma M))^{2}}{8|\gamma|}\\ &\qquad+\frac{\eta A^{2}(\gamma)}{8n}+\frac{(\mathrm{KL}(\rho_{h}\|Q_{h})+\log(1/\delta))}{\eta},\end{split}$ (127) where $A(\gamma)=\exp(|\gamma|M)$, $Q_{h}$ and $\rho_{h}$ are prior and posterior distributions over the hypothesis space, respectively.